### v4

@@ 1. Introduction
We say that a nondeterministic automaton ${\mathcal{A}}$ is almost-DBP if ${\mathcal{A}}$ embodies a deterministic automaton ${\mathcal{A}}^{\prime}$ such that the probability of a random word to be in $L({\mathcal{A}})\setminus L({\mathcal{A}}^{\prime})$ is 00. Clearly, if ${\mathcal{A}}$ is DBP, then it is almost-DBP. Recall that DBPness is associated with the ability of an online algorithm to perform as good as an offline one. A typical analysis of the performance of an online algorithm compares its performance with that of an offline algorithm. The notion of almost-DBPness captures cases where the online algorithm performs, with probability $1$, as good as the offline algorithm. From the point of view of formal language theory, almost-DBPness captures the ability to approximate from below the language of automata that need not be DBP. We study the almost-DBPness of HD and SD automata. Our results imply that the nondeterminism hierarchy “almost collapses”: We show that for Büchi (and hence also weak) automata, semantic determinism implies almost-DBPness, thus every SD-NBW is almost-DBP. Then, for co-Büchi automata, semantic determinism is not enough, and we need HDness. Thus, there is an SD-NCW that is not almost-DBP, yet all HD-NCWs are almost-DBP.
@@ 2.1. Automata
[…]
@@ 3. The Syntactic and Semantic Hierarchies
[Semantic Hierarchy] For $X\in\{B,W\}$, we have that DXW $=$ DBP-NXW $=$ HD-NXW $=$ SD-NXW $<$ NXW.
@@
In [KSV06, KS15], the authors suggest variants of the subset construction that determinize HD-NBWs. As we argue below, the construction in [KS15] is correct also when applied to SD-NBWs. Moreover, it preserves weakness. Thus, DBW$=$SD-NBW and DWW$=$SD-NWW. Also, the last inequality follows from the fact DBW$<$NBW and DWW$<$NWW [Lan69].
Given an NBW ${\mathcal{A}}=\left\langle\Sigma,Q,q_{0},\delta,\alpha\right\rangle$, the DBW generated in [KS15] [^4] The construction in [KS15] assumes automata with transition-based acceptance, and (regardless of this) is slightly different: when $\alpha$ is visited, ${\mathcal{A}}^{\prime}$ continues with a single state from the set of successors. However, the key point remains the same: ${\mathcal{A}}$ being SD enables ${\mathcal{A}}^{\prime}$ to maintain only subsets of states, rather than Safra trees, which makes determinization much easier. is ${\mathcal{A}}^{\prime}=\left\langle\Sigma,Q^{\prime},q^{\prime}_{0},\delta^{\prime},\alpha^{\prime}\right\rangle$, where $Q^{\prime}=2^{Q}$, $q^{\prime}_{0}=\{q_{0}\}$, $\alpha^{\prime}=\{S\in 2^{Q}:S\subseteq\alpha\}$, and the transition function $\delta^{\prime}$ is defined for every subset $S\in 2^{Q}$ and letter $\sigma\in\Sigma$ as follows. If $\delta(S,\sigma)\cap\alpha=\emptyset$, then $\delta^{\prime}(S,\sigma)=\delta(S,\sigma)$. Otherwise, namely if $\delta(S,\sigma)\cap\alpha\neq\emptyset$, then $\delta^{\prime}(S,\sigma)=\delta(S,\sigma)\cap\alpha$.
@@ 4. Deciding the Nondeterminism Level of an Automaton
Consider a SAT instance $\varphi$ over the variable set $X=\{x_{1},\ldots,x_{n}\}$ and with $m\geq 1$ clauses $C=\{c_{1},\ldots,c_{m}\}$. For $n\geq 1$, let $[n]=\{1,2,\ldots,n\}$. For a variable $x_{k}\in X$, let $C^{0}_{k}\subseteq C$ be the set of clauses in which $x_{k}$ appears negatively, and let $C^{1}_{k}\subseteq C$ be the set of clauses in which $x_{k}$ appears positively. For example, if $c_{1}=x_{1}\lor\lnot x_{2}\lor x_{3}$, then $c_{1}$ is in $C^{1}_{1}$, $C_{2}^{0}$, and $C^{1}_{3}$. Assume that all clauses depend on at least two different variables (that is, no clause is a tautology or forces an assignment to a single variable). Let $\Sigma_{n,m}=X\cup C$, and let
@@
The upper bound follows by Lemma 18 in [KM18] where the authors prove that checking if an automaton is DBP is in NP already for general NCWs. For the lower bound, we analyze the reduction in [KM18] and prove that the NCW constructed there is HD. The reduction is from the Hamiltonian-cycle problem: given a connected graph $G=\left\langle[n],E\right\rangle$, the reduction outputs an NCW ${\mathcal{A}}_{G}$ over the alphabet $[n]$ that is obtained from $G$ by adding self loops to all vertices, labelling the loop at a vertex $i$ by the letter $i$, and labelling the edges from vertex $i$ to all its neighbours in $G$ by every letter $j\neq i$. Then, the co-Büchi condition requires a run to eventually get stuck at a self-loop [^7] The exact reduction is more complicated and involves an additional letter ##\## that forces each deterministic pruning of ${\mathcal{A}}_{G}$ to proceed to the same neighbour of $i$ upon reading a letter $j\neq i$ from the vertex $i$.. Accordingly, $L({\mathcal{A}}_{G})=[n]^{*}\cdot\bigcup_{i\in[n]}i^{\omega}$.
It is not hard to see that ${\mathcal{A}}_{G}$ is HD. Indeed, an HD strategy can decide to which neighbour of $i$ to proceed with a letter $j\neq i$ by following a cycle $c$ that traverses all the vertices of the graph $G$. Since when we read $j\neq i$ at vertex $i$ we move to a neighbour state, then by following the cycle $c$ upon reading $i^{\omega}$, we eventually reach the vertex $i$ and get stuck at the $i$-labeled loop. Thus, the Hamiltonian-cycle problem is reduced to DBPness of an HD-NCW, and we are done.
@@ 5. Approximated Determinization by Pruning
[…]
@@
Let ${\mathcal{A}}$ be an SD-NBW, and consider the simple stochastic Büchi game $\mathcal{G}_{\mathcal{A}}$. We first show that Eve has a probabilistic strategy to win ${\mathcal{G}}_{\mathcal{A}}$ with probability $1$, even without assuming that ${\mathcal{A}}$ is semantically deterministic. Consider the probabilistic strategy $g$ where from $(q,\sigma)$, Eve picks one of the $\sigma$-successors of $q$ uniformly at random. Note that this strategy is memoryless, and hence the outcome of the game can be thought as a random walk in ${\mathcal{A}}$ that starts at $q_{0}$ and gives positive probabilities to all transitions. Thus, with probability $1$, the run $r_{g}$ is going to reach an ergodic SCC of ${\mathcal{A}}$ and visit all its states. If the run $r_{g}$ reaches an $\alpha$-free ergodic SCC, then $\inf(r_{g})\subseteq Q_{rej}$, and hence $r_{g}$ is then winning for Eve. Otherwise, $r_{g}$ reaches a non $\alpha$-free ergodic SCC, and with probability $1$, it visits all the states in that SCC. Thus, with probability $1$, we have $\inf(r_{g})\cap\alpha\neq 0$, and $r_{g}$ is winning for Eve. Overall, Eve wins ${\mathcal{G}}_{\mathcal{A}}$ with probability $1$ when playing according to $g$.
For a strategy $f$ for Eve, we say that $r_{f}$ is correct if $w_{f}\in L({\mathcal{A}})$ implies that $r_{f}$ is accepting. Note that $w_{f}\notin L({\mathcal{A}})$ always implies that $r_{f}$ is rejecting. We show that if ${\mathcal{A}}$ is SD, then for every almost-sure winning strategy $f$ for Eve, we have that $r_{f}$ is correct with probability $1$. Consider an almost-sure winning strategy $f$ for Eve. Since $f$ is almost-sure winning for Eve, we have that with probability 1, either $r_{f}$ is an accepting run or $\inf(r_{f})\subseteq Q_{rej}$. Thus, it is sufficient to prove that if $\inf(r_{f})\subseteq Q_{rej}$, then $w_{f}\in L({\mathcal{A}})$ with probability 00, since otherwise, almost surely $r_{f}$ is accepting, and in particular is correct.
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@ 6.1. Good-for-MDP automata
Let ${\mathcal{M}}=\left\langle S,s_{0},(A_{s})_{s\in S},\mathrm{Pr}_{\mathcal{M}},\Sigma,\tau\right\rangle$ be an MDP and ${\mathcal{N}}=\left\langle\Sigma,Q,q_{0},\delta,\alpha\right\rangle$ be a nondeterministic automaton. When the player takes actions in ${\mathcal{M}}$ the infinite word $\tau(r)\in\Sigma^{\omega}$ is generated. Assume that, simultaneously, the player also generates a run of ${\mathcal{N}}$. This process can be modeled by an MDP that is a product of ${\mathcal{M}}$ with ${\mathcal{N}}$, and in each step the player takes an action of ${\mathcal{M}}$ and chooses a successor state in ${\mathcal{N}}$ that extends the current run on the new letter drawn by ${\mathcal{M}}$. Formally, we define the MDP ×${\mathcal{M}}\times{\mathcal{N}}=\left\langle S^{\prime},\left\langle s_{0},\bot\right\rangle,A^{\prime},\mathrm{Pr}_{{\mathcal{M}}\times{\mathcal{N}}},\Sigma,\gamma\right\rangle$ that is defined as follows. The set of states of ×${\mathcal{M}}\times{\mathcal{N}}$ is $S^{\prime}=(S\times Q)\cup\left\{\left\langle s_{0},\bot\right\rangle\right\}$, for $\bot\notin Q$. The set of actions $A^{\prime}=A\left\langle s_{0},\bot\right\rangle\cup\bigcup_{\left\langle s,q\right\rangle\in S\times Q}(A_{\left\langle s,q\right\rangle})$ is defined as follows. For all $\left\langle s,q\right\rangle\in S\times Q$, the set of actions $A_{\left\langle s,q\right\rangle}$ available from $\left\langle s,q\right\rangle$ consists of all pairs $\left\langle a,q^{\prime}\right\rangle\in A_{s}\times\delta(q,\tau(s))$. In addition, the actions available from $\left\langle s_{0},\bot\right\rangle$ are $A_{\left\langle s_{0},\bot\right\rangle}=A_{s_{0}}\times\left\{q_{0}\right\}$. Then, for all $\left\langle s,\ast\right\rangle\in S^{\prime}$ and $\left\langle a,q^{\prime}\right\rangle\in A_{\left\langle s,\ast\right\rangle}$, we have $\mathrm{Pr}_{{\mathcal{M}}\times{\mathcal{N}}}(\left\langle s,\ast\right\rangle,\left\langle a,q^{\prime}\right\rangle,\left\langle s^{\prime},q^{\prime}\right\rangle)=\mathrm{Pr}_{{\mathcal{M}}}(s,a,s^{\prime})$ and $\gamma(s,\ast)=\tau(s)$. Given a finite or infinite run $r\in\Pi({\mathcal{M}}\times{\mathcal{N}})\cup\Omega({\mathcal{M}}\times{\mathcal{N}})$, $r=\left\langle\left\langle s_{0},\bot\right\rangle,\left\langle s_{1},q_{0}\right\rangle,\left\langle s_{2},q_{1}\right\rangle,\ldots\right\rangle$, we define $\pi_{\mathcal{N}}(r)=\left\langle q_{0},q_{1},q_{2},\ldots\right\rangle$, and $\pi_{\mathcal{M}}(r)=\left\langle s_{0},s_{1},s_{2},\ldots\right\rangle$ to be the ${\mathcal{N}}$-run and ${\mathcal{M}}$-run that correspond to $r$ respectively. Conversely, for $r_{\mathcal{M}}=\left\langle s_{0},s_{1},s_{2},\ldots\right\rangle\in S^{*}\cup S^{\omega}$ and $r_{\mathcal{N}}=\left\langle q_{0},q_{1},q_{2},\ldots\right\rangle\in(Q_{\mathcal{N}})^{*}\cup(Q_{\mathcal{N}})^{\omega}$ of appropriate lengths, let $r_{\mathcal{M}}\oplus r_{\mathcal{N}}=\left\langle\left\langle s_{0},\bot\right\rangle,\left\langle s_{1},q_{0}\right\rangle,\left\langle s_{2},q_{1}\right\rangle,\ldots\right\rangle$. Notice that for all $r_{\mathcal{M}}\oplus r_{\mathcal{N}}\in\Pi({\mathcal{M}}\times{\mathcal{N}})\cup\Omega({\mathcal{M}}\times{\mathcal{N}})$, we have $\pi_{\mathcal{N}}(r_{\mathcal{M}}\oplus r_{\mathcal{N}})=r_{\mathcal{N}}$ and $\pi_{\mathcal{M}}(r_{\mathcal{M}}\oplus r_{\mathcal{N}})=r_{\mathcal{M}}$.
@@
Maximizing the syntactic satisfaction can be expressed as a simple stochastic parity game that is played on top of the MDP ×${\mathcal{M}}\times{\mathcal{N}}$, where the winning objective is inherited from ${\mathcal{N}}$. In  [CH03], the authors prove that simple stochastic parity games admit optimal pure memoryless strategies. Thus, the supremum in the definition of $\mathrm{PSyn}^{\mathcal{N}}_{{\mathcal{M}}}$ is attained by a pure memoryless strategy. Note that $\mathrm{PSyn}^{\mathcal{N}}_{{\mathcal{M}}}$ is evaluated with respect to the MDP ×${\mathcal{M}}\times{\mathcal{N}}$, and thus by pure memoryless we mean that there is a strategy $\mu^{*}_{\mathcal{N}}:S\times Q\cup\left\{\left\langle s_{0},\bot\right\rangle\right\}\rightarrow A\times Q$ such that $\mathrm{PSyn}^{\mathcal{N}}_{{\mathcal{M}}}(\mu^{*}_{\mathcal{N}})=\mathrm{PSyn}^{\mathcal{N}}_{{\mathcal{M}}}$.
@@
Similarly, the semantic satisfaction can be expressed as a simple stochastic parity game that augments ×${\mathcal{M}}\times{\mathcal{N}}$ with a deterministic automaton ${\mathcal{D}}$ for $L({\mathcal{N}})$. Thus, the supremum in the definition of $\mathrm{PSem}^{\mathcal{N}}_{{\mathcal{M}}}$ is attained by a pure finite-memory strategy. Indeed, the considerations are as in the syntactic case, except that here, the current state of the deterministic automaton ${\mathcal{D}}$ is used as a memory for the player.

### v5

@@ 1. Introduction
We say that a nondeterministic automaton $\mathcal{A}$} is almost-DBP if $\mathcal{A}$} embodies a deterministic automaton $\mathcal{A}$$^{\prime}$ such that the probability of a random word to be in $L({\mathcal{A}})\setminus L({\mathcal{A}}^{\prime})$ is 0. Clearly, if $\mathcal{A}$} is DBP, then it is almost-DBP. Recall that DBPness is associated with the ability of an online algorithm to perform as good as an offline one. A typical analysis of the performance of an online algorithm compares its performance with that of an offline algorithm. The notion of almost-DBPness captures cases where the online algorithm performs, with probability $1$, as good as the offline algorithm. From the point of view of formal language theory, almost-DBPness captures the ability to approximate from below the language of automata that need not be DBP. We study the almost-DBPness of HD and SD automata. Our results imply that the nondeterminism hierarchy “almost collapses”: We show that for Büchi (and hence also weak) automata, semantic determinism implies almost-DBPness, thus every SD-NBW is almost-DBP. Then, for co-Büchi automata, semantic determinism is not enough, and we need HDness. Thus, there is an SD-NCW that is not almost-DBP, yet all HD-NCWs are almost-DBP.
@@ 2.1. Automata
The acceptance condition $\alpha$ determines which runs are “good”. We consider here the Büchi and co-Büchi acceptance conditions, where $\alpha\subseteq Q$ is a subset of states. We use the terms $\alpha$-states and $\bar{\alpha}$-states to refer to states in $\alpha$ and in $Q\setminus\alpha$, respectively. For a run $r$, let ${\it inf}(r)\subseteq Q$ be the set of states that $r$ traverses infinitely often. Thus, ${\it inf}(r)=\{q\in Q:q=r_{i}\text{ for infinitely many $i$'s}\}$. A run $r$ of a Büchi automaton is accepting iff it visits states in $\alpha$ infinitely often, thus ${\it inf}(r)\cap\alpha\neq\emptyset$. Dually, a run $r$ of a co-Büchi automaton is accepting iff it visits states in $\alpha$ only finitely often, thus ${\it inf}(r)\cap\alpha=\emptyset$. A run that is not accepting is rejecting. Note that as $\mathcal{A}$} is nondeterministic, it may have several runs on a word $w$. The word $w$ is accepted by $\mathcal{A}$} if there is an accepting run of $\mathcal{A}$} on $w$. The language of $\mathcal{A}$}, denoted $L({\mathcal{A}})$, is the set of words that $\mathcal{A}$} accepts. Two automata are equivalent if their languages are equivalent.
@@ 3. The Syntactic and Semantic Hierarchies
[Semantic Hierarchy] For $X\in\{B,W\}$, we have that DXW == DBP-NXW == HD-NXW == SD-NXW << NXW.
@@
In [KSV06, KS15], the authors suggest variants of the subset construction that determinize HD-NBWs. As we argue below, the construction in [KS15] is correct also when applied to SD-NBWs. Moreover, it preserves weakness. Thus, DBW==SD-NBW and DWW==SD-NWW. Also, the last inequality follows from the fact DBW<<NBW and DWW<<NWW [Lan69].
Given an NBW ${\mathcal{A}}=\left\langle\Sigma,Q,q_{0},\delta,\alpha\right\rangle$, the DBW ${\mathcal{A}}^{\prime}=\left\langle\Sigma,Q^{\prime},q^{\prime}_{0},\delta^{\prime},\alpha^{\prime}\right\rangle$ generated in [KS15] [^4] The construction in [KS15] assumes automata with transition-based acceptance, and (regardless of this) is slightly different: when $\alpha$ is visited, $\mathcal{A}$$^{\prime}$ continues with a single state from the set of successors. However, the key point remains the same: $\mathcal{A}$} being SD enables $\mathcal{A}$$^{\prime}$ to maintain only subsets of states, rather than Safra trees, which makes determinization much easier. is such that $Q^{\prime}=2^{Q}$, $q^{\prime}_{0}=\{q_{0}\}$, $\alpha^{\prime}=\{S\in 2^{Q}:S\subseteq\alpha\}$, and the transition function $\delta^{\prime}$ is defined for every subset $S\in 2^{Q}$ and letter $\sigma\in\Sigma$ as follows. If $\delta(S,\sigma)\cap\alpha=\emptyset$, then $\delta^{\prime}(S,\sigma)=\delta(S,\sigma)$. Otherwise, namely if $\delta(S,\sigma)\cap\alpha\neq\emptyset$, then $\delta^{\prime}(S,\sigma)=\delta(S,\sigma)\cap\alpha$.
@@ 4. Deciding the Nondeterminism Level of an Automaton
Consider a SAT instance $\varphi$ over the variable set $X=\{x_{1},\ldots,x_{n}\}$ and with $m\geq 1$ clauses $C=\{c_{1},\ldots,c_{m}\}$. For $n\geq 1$, let $n]=\{1,2,\ldots,n\}$. For a variable $x_{k}\in X$, let $C^{0}_{k}\subseteq C$ be the set of clauses in which $x_{k}$ appears negatively, and let $C^{1}_{k}\subseteq C$ be the set of clauses in which $x_{k}$ appears positively. For example, if $c_{1}=x_{1}\lor\lnot x_{2}\lor x_{3}$, then $c_{1}$ is in $C^{1}_{1}$, $C_{2}^{0}$, and $C^{1}_{3}$. Assume that all clauses depend on at least two different variables (that is, no clause is a tautology or forces an assignment to a single variable). Let $\Sigma_{n,m}=X\cup C$, and let
@@
The upper bound follows by Lemma 18 in [KM18] where the authors prove that checking if an automaton is DBP is in NP already for general NCWs. For the lower bound, we analyze the reduction in [KM18] and prove that the NCW constructed there is HD. The reduction is from the Hamiltonian-cycle problem: given a connected graph $G=\left\langle[n],E\right\rangle$, the reduction outputs an NCW ${\mathcal{A}}_{G}$ over the alphabet $[n]$ that is obtained from $G$ by adding self loops to all vertices, labelling the loop at a vertex $i$ by the letter $i$, and labelling the edges from vertex $i$ to all its neighbors in $G$ by every letter $j\neq i$. Then, the co-Büchi condition requires a run to eventually get stuck at a self-loop [^7] The exact reduction is more complicated and involves an additional letter #\# that forces each deterministic pruning of ${\mathcal{A}}_{G}$ to proceed to the same neighbor of $i$ upon reading a letter $j\neq i$ from the vertex $i$.. Accordingly, $L({\mathcal{A}}_{G})=[n]^{*}\cdot\bigcup_{i\in[n]}i^{\omega}$.
It is not hard to see that ${\mathcal{A}}_{G}$ is HD. Indeed, an HD strategy can decide to which neighbor of $i$ to proceed with a letter $j\neq i$ by following a cycle $c$ that traverses all the vertices of the graph $G$. Since when we read $j\neq i$ at vertex $i$ we move to a neighbor state, then by following the cycle $c$ upon reading $i^{\omega}$, we eventually reach the vertex $i$ and get stuck at the $i$-labeled loop. Thus, the Hamiltonian-cycle problem is reduced to DBPness of an HD-NCW, and we are done.
@@ 5. Approximated Determinization by Pruning
We start with the positive result about Büchi automata. Consider an NBW ${\mathcal{A}}=\left\langle\Sigma,Q,q_{0},\delta,\alpha\right\rangle$. We define a simple stochastic Büchi game $\mathcal{G}_{\mathcal{A}}$ as follows. [^9] In [CJH03] these games are called simple $1\frac{1}{2}$ -player games with Büchi winning objectives and almost-sure winning criterion. The game is played between Random and Eve. The set of positions of Random are $Q$, and the set of positions of Eve are $Q\times\Sigma$. The game starts from position $q_{0}$. A round in the game starts at some position $q\in Q$ and proceeds as follows.
@@
Let $\mathcal{A}$} be an SD-NBW, and consider the simple stochastic Büchi game $\mathcal{G}_{\mathcal{A}}$. We first show that Eve has a probabilistic strategy to win ${\mathcal{G}}_{\mathcal{A}}$ with probability $1$, even without assuming that $\mathcal{A}$} is semantically deterministic. Consider the probabilistic strategy $g$ where from $(q,\sigma)$, Eve picks one of the $\sigma$-successors of $q$ uniformly at random. Note that this strategy is memoryless, and hence the outcome of the game can be thought as a random walk in $\mathcal{A}$} that starts at $q_{0}$ and gives positive probabilities to all transitions. Thus, with probability $1$, the run $r_{g}$ is going to reach an ergodic SCC of $\mathcal{A}$} and visit all its states. If the run $r_{g}$ reaches an $\alpha$-free ergodic SCC, then $\inf(r_{g})\subseteq Q_{rej}$, and hence $r_{g}$ is then winning for Eve. Otherwise, $r_{g}$ reaches a non $\alpha$-free ergodic SCC, and with probability $1$, it visits all the states in that SCC. Thus, with probability 11, we have $\inf(r_{g})\cap\alpha\neq 0$, and $r_{g}$ is winning for Eve. Overall, Eve wins ${\mathcal{G}}_{\mathcal{A}}$ with probability $1$ when playing according to $g$.
For a strategy $f$ for Eve, we say that $r_{f}$ is correct if $w_{f}\in L({\mathcal{A}})$ implies that $r_{f}$ is accepting. Note that $w_{f}\notin L({\mathcal{A}})$ always implies that $r_{f}$ is rejecting. We show that if $\mathcal{A}$} is SD, then for every almost-sure winning strategy $f$ for Eve, we have that $r_{f}$ is correct with probability $1$. Consider an almost-sure winning strategy $f$ for Eve. Since $f$ is almost-sure winning for Eve, we have that with probability 1, either $r_{f}$ is an accepting run or $\inf(r_{f})\subseteq Q_{rej}$. Thus, it is sufficient to prove that if $\inf(r_{f})\subseteq Q_{rej}$, then $w_{f}\in L({\mathcal{A}})$ with probability 0, since otherwise, almost surely $r_{f}$ is accepting, and in particular is correct.
@@
Consider a language $L\subseteq\Sigma^{\omega}$ of infinite words. We say that a finite word $x\in\Sigma^{*}$ is a good prefix for $L$ if $x\cdot\Sigma^{\omega}\subseteq L$. Then, $L$ is a co-safety language if every word in $L$ has a good prefix [AS87]. Let $\mbox{\it co-safe}(L)=\{x\cdot w\in\Sigma^{\omega}:\text{$x$ is a good prefix of }L\}$. Clearly, $\mbox{\it co-safe}(L)\subseteq L$. The other direction is not necessarily true. For example, if $L\subseteq\{a,b\}^{\omega}$ is the set of all words with infinitely many $a$’s, then $\mbox{\it co-safe}(L)=\emptyset$. In fact, $\mbox{\it co-safe}(L)=L$ iff $L$ is a co-safety language. As we show now, when $L$ is NCW-recognizable, we can relate $L$ and $\mbox{\it co-safe}(L)$ as follows.
@@
By Lemma 12, pruning an NCW in a way that would make it recognize $\mbox{\it co-safe}(L({\mathcal{A}}))$ results in a DCW that approximates $\mathcal{A}$}, and thus witnesses that $\mathcal{A}$} is almost-DBP. We now show that for HD-NCWs, such a pruning is possible, and conclude that HD-NCWs are almost-DBP.
@@
In order to see that $\mathcal{A}$$^{\prime}$ is HD and that $\mbox{\it co-safe}(L({\mathcal{A}}))\subseteq L({\mathcal{A}}^{\prime})$, consider an HD strategy $f$ of $\mathcal{A}$}. We claim that $f$ is also an HD strategy for $\mathcal{A}$$^{\prime}$, and for all words $w\in\mbox{\it co-safe}(L({\mathcal{A}}))$, the run $r$ that $f$ generates on $w$ eventually visits only states $q\notin\alpha^{\prime}$. Since $\mbox{\it co-safe}(L({\mathcal{A}}))\subseteq L({\mathcal{A}})$, we know that $\inf(r)\cap\alpha=\emptyset$. It is left to prove that $r$ visits only finitely many states $q\in Q$ with $L({\mathcal{A}}^{q})\neq\Sigma^{\omega}$.
@@
So, $\mathcal{A}$$^{\prime}$ is an HD-NCW with $L({\mathcal{A}}^{\prime})=\mbox{\it co-safe}(L({\mathcal{A}}))$. Since $\mbox{\it co-safe}(L({\mathcal{A}}))$ is co-safe, it is DWW-recognizable [Sis94]. By [BKS17], HD-NCWs whose language is DWW-realizable are DBP. Let $\delta^{\prime}$ be the restriction of $\delta$ to a deterministic transition function such that ${\mathcal{D}}^{\prime}=\left\langle\Sigma,Q,q_{0},\delta^{\prime},\alpha^{\prime}\right\rangle$ is a DCW with $L({\mathcal{D}}^{\prime})=L({\mathcal{A}}^{\prime})=\mbox{\it co-safe}(L({\mathcal{A}}))$. Consider now the DCW ${\mathcal{D}}=\left\langle\Sigma,Q,q_{0},\delta^{\prime},\alpha\right\rangle$ that is obtained from $\mathcal{D}$$^{\prime}$ by replacing $\alpha^{\prime}$ with $\alpha$.
It is clear that $\mathcal{D}$} is a pruning of $\mathcal{A}$}. Note that, $\alpha\subseteq\alpha^{\prime}$, and hence $\mbox{\it co-safe}(L({\mathcal{A}}))=L({\mathcal{D}}^{\prime})\subseteq L({\mathcal{D}})$. That is, $\mathcal{D}$} is a pruning of $\mathcal{A}$} that approximates $L({\mathcal{A}})$ up to a negligible set, and $\mathcal{A}$} is almost-DBP.
@@ 6.1. Good-for-MDP automata
Let ${\mathcal{M}}=\left\langle S,s_{0},(A_{s})_{s\in S},\mathrm{Pr}_{\mathcal{M}},\Sigma,\tau\right\rangle$ be an MDP and ${\mathcal{N}}=\left\langle\Sigma,Q,q_{0},\delta,\alpha\right\rangle$ be a nondeterministic automaton. When the player takes actions in $\mathcal{M}$} the infinite word $\tau(r)\in\Sigma^{\omega}$ is generated. Assume that, simultaneously, the player also generates a run of $\mathcal{N}$}. This process can be modeled by an MDP that is a product of $\mathcal{M}$} with $\mathcal{N}$}, and in each step the player takes an action of $\mathcal{M}$} and chooses a successor state in $\mathcal{N}$} that extends the current run on the new letter drawn by $\mathcal{M}$}. Formally, we define the MDP ${\mathcal{M}}\times{\mathcal{N}}=\left\langle S^{\prime},\left\langle s_{0},\bot\right\rangle,A^{\prime},\mathrm{Pr}_{{\mathcal{M}}\times{\mathcal{N}}},\Sigma,\gamma\right\rangle$ that is defined as follows. The set of states of ${\mathcal{M}}\times{\mathcal{N}}$ is $S^{\prime}=(S\times Q)\cup\left\{\left\langle s_{0},\bot\right\rangle\right\}$, for $\bot\notin Q$. The set of actions $A^{\prime}=A\left\langle s_{0},\bot\right\rangle\cup\bigcup_{\left\langle s,q\right\rangle\in S\times Q}(A_{\left\langle s,q\right\rangle})$ is defined as follows. For all $\left\langle s,q\right\rangle\in S\times Q$, the set of actions $A_{\left\langle s,q\right\rangle}$ available from $\left\langle s,q\right\rangle$ consists of all pairs $\left\langle a,q^{\prime}\right\rangle\in A_{s}\times\delta(q,\tau(s))$. In addition, the actions available from $\left\langle s_{0},\bot\right\rangle$ are $A_{\left\langle s_{0},\bot\right\rangle}=A_{s_{0}}\times\left\{q_{0}\right\}$. Then, for all $\left\langle s,\ast\right\rangle\in S^{\prime}$ and $\left\langle a,q^{\prime}\right\rangle\in A_{\left\langle s,\ast\right\rangle}$, we have $\mathrm{Pr}_{{\mathcal{M}}\times{\mathcal{N}}}(\left\langle s,\ast\right\rangle,\left\langle a,q^{\prime}\right\rangle,\left\langle s^{\prime},q^{\prime}\right\rangle)=\mathrm{Pr}_{{\mathcal{M}}}(s,a,s^{\prime})$ and $\gamma(s,\ast)=\tau(s)$. Given a finite or infinite run $r\in\Pi({\mathcal{M}}\times{\mathcal{N}})\cup\Omega({\mathcal{M}}\times{\mathcal{N}})$, $r=\left\langle\left\langle s_{0},\bot\right\rangle,\left\langle s_{1},q_{0}\right\rangle,\left\langle s_{2},q_{1}\right\rangle,\ldots\right\rangle$, we define $\pi_{\mathcal{N}}(r)=\left\langle q_{0},q_{1},q_{2},\ldots\right\rangle$, and $\pi_{\mathcal{M}}(r)=\left\langle s_{0},s_{1},s_{2},\ldots\right\rangle$ to be the $\mathcal{N}$}-run and $\mathcal{M}$}-run that correspond to $r$ respectively. Conversely, for $r_{\mathcal{M}}=\left\langle s_{0},s_{1},s_{2},\ldots\right\rangle\in S^{*}\cup S^{\omega}$ and $r_{\mathcal{N}}=\left\langle q_{0},q_{1},q_{2},\ldots\right\rangle\in(Q_{\mathcal{N}})^{*}\cup(Q_{\mathcal{N}})^{\omega}$ of appropriate lengths, let $r_{\mathcal{M}}\oplus r_{\mathcal{N}}=\left\langle\left\langle s_{0},\bot\right\rangle,\left\langle s_{1},q_{0}\right\rangle,\left\langle s_{2},q_{1}\right\rangle,\ldots\right\rangle$. Notice that for all $r_{\mathcal{M}}\oplus r_{\mathcal{N}}\in\Pi({\mathcal{M}}\times{\mathcal{N}})\cup\Omega({\mathcal{M}}\times{\mathcal{N}})$, we have $\pi_{\mathcal{N}}(r_{\mathcal{M}}\oplus r_{\mathcal{N}})=r_{\mathcal{N}}$ and $\pi_{\mathcal{M}}(r_{\mathcal{M}}\oplus r_{\mathcal{N}})=r_{\mathcal{M}}$.
@@
Maximizing the syntactic satisfaction can be expressed as a simple stochastic parity game that is played on top of the MDP ${\mathcal{M}}\times{\mathcal{N}}$, where the winning objective is inherited from $\mathcal{N}$}. In  [CH03], the authors prove that simple stochastic parity games admit optimal pure memoryless strategies. Thus, the supremum in the definition of $\mathrm{PSyn}^{\mathcal{N}}_{{\mathcal{M}}}$ is attained by a pure memoryless strategy. Note that $\mathrm{PSyn}^{\mathcal{N}}_{{\mathcal{M}}}$ is evaluated with respect to the MDP ${\mathcal{M}}\times{\mathcal{N}}$, and thus by pure memoryless we mean that there is a strategy $\mu^{*}_{\mathcal{N}}:S\times Q\cup\left\{\left\langle s_{0},\bot\right\rangle\right\}\rightarrow A\times Q$ such that $\mathrm{PSyn}^{\mathcal{N}}_{{\mathcal{M}}}(\mu^{*}_{\mathcal{N}})=\mathrm{PSyn}^{\mathcal{N}}_{{\mathcal{M}}}$.
@@
Similarly, the semantic satisfaction can be expressed as a simple stochastic parity game that augments ${\mathcal{M}}\times{\mathcal{N}}$ with a deterministic automaton $\mathcal{D}$} for $L({\mathcal{N}})$. Thus, the supremum in the definition of $\mathrm{PSem}^{\mathcal{N}}_{{\mathcal{M}}}$ is attained by a pure finite-memory strategy. Indeed, the considerations are as in the syntactic case, except that here, the current state of the deterministic automaton $\mathcal{D}$} is used as a memory for the player.